solve the equation
2x^3-3x^2-5x=0
can u show me the steps

Answer by sudhanshu_kmr(1152)


2x^3-3x^2-5x=0
=> x(2x^2-3x-5)=0
=> x( 2x^2-5x +2x -5)=0
=> x(x(2x-5)+1(2x-5))=0
=> x(x+1)(2x-5)=0
x = 0, -1 or 5/2